OpenStudy (anonymous):

Kenny owned so many baseball cards that he sold 82 of them after selling the cards he still had 159 left write and solve an equation to find the number of baseball cards b Kenny had originally.

OpenStudy (anonymous):

b - 82 = 159 and you try to solve it, if you can't I will try and help you again.

OpenStudy (anonymous):

159 + 82 =

OpenStudy (anonymous):

b + 159 = 82 241 cards

OpenStudy (anonymous):

241 is the answer because if he gave 82 away and then had left 159 we do... 82+159= 241

OpenStudy (anonymous):

I will give medal